OpenStudy (love_to_love_you):

Please hellppp. Medal and fan. Identify the slope of the line that passes through the points (5, 2) and (-4, 6) show steps.

OpenStudy (kl0723):

y2-y1/x2-x1

OpenStudy (kl0723):

choose x1=5 x2=2 y1=-4 and y2=6

OpenStudy (texaschic101):

you will need to use the slope formula slope = (y2 - y1) / (x2 - x1) (5,2)...x1 = 5 and y1 = 2 (-4,6)..x2 = -4 and y2 = 6 now we sub slope = (6 - 2) / (-4 - 5) can you finish ?